OpenStudy (anonymous):

DIFFERENTIAL EQUATIONS: Given the following nonlinear system, graph the direction of the nullclines and final phase portrait

OpenStudy (anonymous):

I'm having trouble deciding the direction to graph the nullclines when they are not in equilibrium. Can you please help? dx/dt = x(10-x-y) dy/dt=y(30-2x-y)